Got all the multiple choice ones correct just need help with the wrong answers
Test the claim that the proportion of people who own cats is larger than 80% at the 0.10 significance level. The null and alternative hypothesis would be Ho: ? < 0.8 Ho: ? > 0.8 Ho:p > 0.8 H! : ? > 0.8 H1 : ? < 0.8 H1 : p < 0.8 The test is: right-tailed left-tailed two-tailed Based on a sample of 600 people, 84% owned cats (to2 The p-value is: C decimals) Based on this we: OReject the null hypothesis OFail to reject the null hypothesis
media%2Fd21%2Fd21e1139-7392-4c1a-8800-a1
Answer #1

Ans:

1)

Test statistic:

z=(0.84-0.8)/sqrt(0.8*(1-0.8)/600)

z=2.45

2)

point estimate=73/180=0.406

90% confidence interval for true proportion

=0.406+/-1.645*sqrt(0.406*(1-0.406)/180)

=0.406+/-0.060

=(0.346, 0.466)